# FairSteer: Inference Time Debiasing for LLMs with Dynamic Activation Steering

### v1

@@ Abstract
Large language models (LLMs) are prone to capturing biases from training corpus, leading to potential negative social impacts. Existing prompt-based debiasing methods exhibit instability due to their sensitivity to prompt changes, while fine-tuning-based techniques incur substantial computational overhead and catastrophic forgetting. In this paper, we propose FairSteer, a novel inference-time debiasing framework without requiring customized prompt design or model retraining. Motivated by the linear representation hypothesis, our preliminary investigation demonstrates that fairness-related features can be encoded into separable directions in the hidden activation space. FairSteer operates in three steps: biased activation detection, debiasing steering vector (DSV) computation, and dynamic activation steering. Specifically, it first trains a lightweight linear classifier to detect bias signatures in activations, and then computes DSVs as intervention directions derived from small contrastive prompt pairs. Subsequently, it performs debiasing by adjusting activations with DSVs in the inference stage. Comprehensive evaluation with six LLMs demonstrates the superiority of FairSteer across question-answering, counterfactual input evaluation and open-ended text generation tasks. Code will be released.
@@
Yichen Li1thanks: Equal contribution., Zhiting Fan [^1] footnotemark: 1, Ruizhe Chen1, Xiaotang Gai1, Luqi Gong2, Yan Zhang1, Zuozhu Liu1thanks: Correspondence author.
1Zhejiang University 2Research Center for Space Computing System, Zhejiang Lab
@@ 1 Introduction
Therefore, the question arises: can we mitigate bias in LLMs without retraining, instead deferring the debiasing process to the inference stage? This approach, which we categorize as inference-time debiasing, is simpler and more practical than in-training debiasing, as it avoids the need for complex training procedures or extensive computational resources. Recent research on inference-time debiasing focuses on modifying decoding strategies (Saunders et al., 2021; Sheng et al., 2020; Meade et al., 2023; Lu et al., 2020b) by suppressing biased tokens. However, such strategies usually reduce the diversity of outputs and degrade the performance of LLMs on non-debiasing tasks.
@@
We conduct comprehensive experiments over six LLMs (Llama, Vicuna, Mistral) and four popular datasets (BBQ (Parrish et al., 2021), UNQOVER (Li et al., 2020), CrowS-Pairs (Nangia et al., 2020) and CEB (Wang et al., 2024)), with tasks across question answering, counterfactual input evaluation, and open-ended text generation. The results demonstrate FairSteer’s superior debiasing performance across different evaluation metrics, while nearly preserving the original performance of LLMs on MMLU/ARC/OBQA knowledge tasks. Extensive analysis, ablation studies, and case studies further highlight the effectiveness of BAD and DAS in FairSteer, suggesting the great potential of inference-time debiasing. Code and data will be publicly released.
@@ 3.1 Biased Activation Detection
where $\lambda$ is the regularization parameter. Implementation details can be found at Appendix B.2.
@@ 4 Experiments
Figure 6: Accuracy across different categories of BBQ for Llama2-13B, comparing the Base Model, CAL, and Ours. Full results are provided in Appendix E.
@@ 4.2 Datasets
Datasets for Biased Activation Detection. We construct the dataset by combining 58,492 examples from BBQ and 10,266 examples from MMLU (Hendrycks et al., 2020). This mixture prevents classifier overfitting to domain-specific artifacts while maintaining discrimination capability across different bias categories. The dataset is split into training and validation sets in a 4:1 ratio, and the classifier is fitted on the training set. Further details are provided in Appendix B.1.
Datasets for DSV Computation. We construct contrast prompt pairs from BBQ. We select BBQ due to its broad coverage of social biases, including nine categories and two intersectional biases. We sample 10 examples from each category, collecting a total of 110 examples to serve as the dataset for DSV computation. Further details and the impact of dataset size are discussed in Appendix C.
@@ 4.3 Evaluation Tasks
Question-Answering. We conduct our experiments on BBQ and UNQOVER. BBQ contains 58,492 questions across nine categories, while UNQOVER includes 40,000 questions across four categories. In both datasets, we use accuracy as the evaluation metric. Additionally, for BBQ, to quantify the extent to which a model systematically provides biased responses, we calculate bias scores separately for ambiguous and disambiguated contexts as defined by Parrish et al. (2021). A detailed description of the bias score calculation is provided in Appendix D.1.
@@
Open-ended text generation. We use CEB and focus on the continuation task related to stereotyping. This dataset is derived from the HolisticBias (Smith et al., 2022) and includes 400 examples across four social groups. We use different bias metrics provided by Dhamala et al. (2021) and Wang et al. (2024) to capture biases from multiple perspectives: (1) Sentiment. We use the VADER (Hutto and Gilbert, 2014) sentiment analyzer to evaluate the sentiments conveyed in the texts generated by LLMs. (2) Toxicity. We use unitaryai/detoxify [^1] https://github.com/unitaryai/detoxify library to measure the toxicity. (3) Regard. Sheng et al. (2019) introduced regard as an alternative to sentiment, measuring the social perception of a demographic in text. Contrary to toxicity, higher scores for regard and sentiment indicate positive connotation. (4) Bias score. we use GPT-4 (Achiam et al., 2023) to assess the degree of stereotyping in the generated outputs, with scores ranging from 0 (nonbiased) to 100 (most biased). Specific prompts from Wang et al. (2024) are provided to GPT-4 to guide the scoring process, as detailed in Appendix D.2.
@@ 4.5 Main Results
Figure 7: Case study examples. Further examples are provided in Appendix F.
@@ 5 Related Work
Researchers have proposed various debiasing methods, which can be group into two main paradigms: (1) In-context prompt-based methods (Dong et al., 2023; Gallegos et al., 2024b; Oba et al., 2024; Sun et al., 2024) leverage carefully crafted instructions or contextual prompts to guide LLMs toward unbiased outputs during generation. These approaches avoid modifying model parameters and are easily applicable to black-box models. However, they are sensitive to prompt formulation, struggle with implicit biases, and depends heavily on prompt quality and specific use cases. (2) Fine-tuning-based methods involve constructing rebalanced datasets and retraining models (Lu et al., 2020a; Webster et al., 2020; Zmigrod et al., 2019; Maudslay et al., 2019; Zayed et al., 2023), or using techniques such as projection-based methods (Bolukbasi et al., 2016; Ravfogel et al., 2020; Liang et al., 2020), debiasing specific components (Gaci et al., 2022; Zayed et al., 2024; Liu et al., 2024; Limisiewicz et al., 2023), contrastive learning (Cheng et al., 2021; He et al., 2022; Oh et al., 2022), adversarial learning (Han et al., 2021; Jin et al., 2020), and reinforcement learning (Liu et al., 2021b; Allam, 2024) approaches. Although effective, these methods involve complex and resource-intensive training, requiring diverse debiasing datasets.
In contrast, inference-time debiasing methods are simpler and more practical, as they do not require retraining or altering the model’s architecture. Existing inference-time techniques primarily focus on decoding strategy modification, such as constrained next-token search (Saunders et al., 2021; Sheng et al., 2020; Meade et al., 2023; Lu et al., 2020b) or generating and reranking alternative outputs (Chung et al., 2023; Kim et al., 2022; Liu et al., 2023, 2021a; Hallinan et al., 2022). However, the key challenge with these methods is balancing effective bias mitigation while preserving output diversity (Gallegos et al., 2024a). In contrast, our method enhances inference-time debiasing by revealing that fairness-related features are linearly separable in activation space, which allows for targeted intervention through geometrically interpretable steering vectors.
@@ A.2 Impact of Social Groups on DSV Computation
Appendix B Details of Biased Activation Detection
B.1 Datasets
@@ B.2 Implemental Details
Appendix C Details of DSV Computation
C.1 Datasets
@@
C.2 Impact of Dataset Size and Random Seed on DSV
@@
Appendix D Evaluation Details
D.1 Bias Score Metrics in BBQ
@@ D.2 Prompt for Scoring on CEB
Appendix E Full Results for Llama2-13B on BBQ
@@
Appendix F Additional Case Studies

### v2

@@ Abstract
Large language models (LLMs) are prone to capturing biases from training corpus, leading to potential negative social impacts. Existing prompt-based debiasing methods exhibit instability due to their sensitivity to prompt changes, while fine-tuning-based techniques incur substantial computational overhead and catastrophic forgetting. In this paper, we propose FairSteer, a novel inference-time debiasing framework without requiring customized prompt design or model retraining. Motivated by the linear representation hypothesis, our preliminary investigation demonstrates that fairness-related features can be encoded into separable directions in the hidden activation space. FairSteer operates in three steps: biased activation detection, debiasing steering vector (DSV) computation, and dynamic activation steering. Specifically, it first trains a lightweight linear classifier to detect bias signatures in activations, and then computes DSVs as intervention directions derived from small contrastive prompt pairs. Subsequently, it performs debiasing by adjusting activations with DSVs in the inference stage. Comprehensive evaluation with six LLMs demonstrates the superiority of FairSteer across question-answering, counterfactual input evaluation and open-ended text generation tasks.
@@
Yichen Li1thanks: Equal contribution., Zhiting Fan [^1] footnotemark: 1, Ruizhe Chen1, Xiaotang Gai1, Luqi Gong3, Yan Zhang4, Zuozhu Liu1,2thanks: Correspondence author.
1Zhejiang University 2Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence 3Zhejiang Lab 4ByteDance {yichen.23, zuozhuliu}@intl.zju.edu.cn
@@ 1 Introduction
Therefore, the question arises: can we mitigate bias in LLMs without retraining, instead deferring the debiasing process to the inference stage? This approach, which we categorize as inference-time debiasing, is simpler and more practical than in-training debiasing, as it avoids the need for complex training procedures or extensive computational resources. Recent research on inference-time debiasing focuses on modifying decoding strategies (Saunders et al., 2021; Sheng et al., 2020; Meade et al., 2023; Lu et al., 2020b; Yang et al., 2022; Schick et al., 2021) by suppressing biased tokens. However, such strategies usually reduce the diversity of outputs and degrade the performance of LLMs on non-debiasing tasks.
@@
We conduct comprehensive experiments over six LLMs (Llama, Vicuna, Mistral) and four popular datasets (BBQ (Parrish et al., 2021), UNQOVER (Li et al., 2020), CrowS-Pairs (Nangia et al., 2020) and CEB (Wang et al., 2024)), with tasks across question answering, counterfactual input evaluation, and open-ended text generation. The results demonstrate FairSteer’s superior debiasing performance across different evaluation metrics, while nearly preserving the original performance of LLMs on MMLU/ARC/OBQA knowledge tasks. Extensive analysis, ablation studies, and case studies further highlight the effectiveness of BAD and DAS in FairSteer, suggesting the great potential of inference-time debiasing. The code is available at https://github.com/LiYichen99/FairSteer.
@@ 3.1 Biased Activation Detection
where $\lambda$ is the regularization parameter. Implementation details can be found at Appendix C.2.
@@ 4 Experiments
Figure 6: Accuracy across different categories of BBQ for Llama2-13B, comparing the Base Model, CAL, and Ours. Full results are provided in Appendix F.
@@ 4.2 Datasets
Datasets for Biased Activation Detection. We construct the dataset by combining 58,492 examples from BBQ and 10,266 examples from MMLU (Hendrycks et al., 2020). This mixture prevents classifier overfitting to domain-specific artifacts while maintaining discrimination capability across different bias categories. The dataset is split into training and validation sets in a 4:1 ratio, and the classifier is fitted on the training set. Further details are provided in Appendix C.1.
Datasets for DSV Computation. We construct contrast prompt pairs from BBQ. We select BBQ due to its broad coverage of social biases, including nine categories and two intersectional biases. We sample 10 examples from each category, collecting a total of 110 examples to serve as the dataset for DSV computation. Further details and the impact of dataset size are discussed in Appendix D.
@@ 4.3 Evaluation Tasks
Question-Answering. We conduct our experiments on BBQ and UNQOVER. BBQ contains 58,492 questions across nine categories, while UNQOVER includes 40,000 questions across four categories. In both datasets, we use accuracy as the evaluation metric. Additionally, for BBQ, to quantify the extent to which a model systematically provides biased responses, we calculate bias scores separately for ambiguous and disambiguated contexts as defined by Parrish et al. (2021). A detailed description of the bias score calculation is provided in Appendix E.1.
@@
Open-ended text generation. We use CEB and focus on the continuation task related to stereotyping. This dataset is derived from the HolisticBias (Smith et al., 2022) and includes 400 examples across four social groups. We use different bias metrics provided by Dhamala et al. (2021) and Wang et al. (2024) to capture biases from multiple perspectives: (1) Sentiment. We use the VADER (Hutto and Gilbert, 2014) sentiment analyzer to evaluate the sentiments conveyed in the texts generated by LLMs. (2) Toxicity. We use unitaryai/detoxify [^1] https://github.com/unitaryai/detoxify library to measure the toxicity. (3) Regard. Sheng et al. (2019) introduced regard as an alternative to sentiment, measuring the social perception of a demographic in text. Contrary to toxicity, higher scores for regard and sentiment indicate positive connotation. (4) Bias score. we use GPT-4 (Achiam et al., 2023) to assess the degree of stereotyping in the generated outputs, with scores ranging from 0 (nonbiased) to 100 (most biased). Specific prompts from Wang et al. (2024) are provided to GPT-4 to guide the scoring process, as detailed in Appendix E.2.
@@ 4.5 Main Results
Figure 7: Case study examples. Further examples are provided in Appendix G.
@@ 5 Related Work
Promoting fairness in LLMs is a critical component of alignment, a broader objective that seeks to ensure AI systems adhere to human values Stiennon et al. (2020); Ouyang et al. (2022); Fan et al. (2024a); Chen et al. (2024d, a, 2025). Researchers have proposed various debiasing methods, which can be group into two main paradigms: (1) In-context prompt-based methods (Dong et al., 2023; Gallegos et al., 2024b; Oba et al., 2024; Sun et al., 2024) leverage carefully crafted instructions or contextual prompts to guide LLMs toward unbiased outputs during generation. These approaches avoid modifying model parameters and are easily applicable to black-box models. However, they are sensitive to prompt formulation, struggle with implicit biases, and depends heavily on prompt quality and specific use cases. (2) Fine-tuning-based methods involve constructing rebalanced datasets and retraining models (Lu et al., 2020a; Webster et al., 2020; Zmigrod et al., 2019; Maudslay et al., 2019; Zayed et al., 2023; Chen et al., 2024c), or using techniques such as projection-based methods (Bolukbasi et al., 2016; Ravfogel et al., 2020; Liang et al., 2020), debiasing specific components (Gaci et al., 2022; Zayed et al., 2024; Liu et al., 2024; Limisiewicz et al., 2023; Chen et al., 2024b), contrastive learning (Cheng et al., 2021; He et al., 2022; Oh et al., 2022), adversarial learning (Han et al., 2021; Jin et al., 2020), and reinforcement learning (Liu et al., 2021b; Allam, 2024) approaches. Although effective, these methods involve complex and resource-intensive training, requiring diverse debiasing datasets.
In contrast, inference-time debiasing methods are simpler and more practical, as they do not require retraining or altering the model’s architecture. Existing inference-time techniques primarily focus on decoding strategy modification, such as constrained next-token search (Saunders et al., 2021; Sheng et al., 2020; Meade et al., 2023; Lu et al., 2020b) or generating and reranking alternative outputs (Chung et al., 2023; Kim et al., 2022; Liu et al., 2023, 2021a; Hallinan et al., 2022; Fan et al., 2024b). However, the key challenge with these methods is balancing effective bias mitigation while preserving output diversity (Gallegos et al., 2024a). Furthermore, many of these techniques (Yang et al., 2022; Schick et al., 2021) are tailored to older models such as BERT (Devlin et al., 2019) and GPT-2 (Radford et al., 2019), limiting their applicability to more recent architectures. Our approach advances inference-time debiasing by demonstrating that fairness-related features are linearly separable in activation space, enabling targeted interventions via geometrically interpretable steering vectors.
@@ 6 Conclusion
Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence, and the Zhejiang University-Angelalign Inc. R&D Center for Intelligent Healthcare.
@@ A.2 Impact of Social Groups on DSV Computation
$\alpha$ Acc BS(a) BS(d)
0 48.60 5.86 2.91
0.1 50.65 4.69 2.72
0.5 64.98 0.96 1.70
1 74.02 -0.82 0.84
1.5 73.99 -0.85 0.82
2 70.87 -0.87 0.81
Table 5: Effect of varying intervention strength $\alpha$ on Llama2-13B on BBQ.
Appendix B Effect of Varying Intervention Strength
In our method, we apply a fixed intervention strength $\alpha$ of DSV, when the bias is detected ($y<0.5$). While simple, this heuristic does not account for the magnitude of bias in activations, which may lead to mildly biased activations and over-steering risk.
Therefore, we conduct an experiment to evaluate the effect of varying (intervention strength) on Llama2-13B’s performance for the BBQ task. As shown in Table 5, increasing marginally reduces BS(a) further but slightly degrades accuracy, suggesting diminishing returns and minor over-steering. While $\alpha=1$ works well globally, we find that local over-steering might occur for mildly biased activations. To mitigate this, we propose an adaptive scaling of the DSV based on the severity of the bias in the future work. Specifically, we can use the classifier’s confidence score $p(y=0)$ to modulate $\alpha$: reduce the intervention strength for borderline cases ($p(y=0)\approx 0.5$), and apply stronger corrections for instances of clear bias ($p(y=0)\gg 0.5$).
Appendix C Details of Biased Activation Detection
C.1 Datasets
@@ C.2 Implemental Details
Appendix D Details of DSV Computation
D.1 Datasets
@@
D.2 Impact of Dataset Size and Random Seed on DSV
@@
Appendix E Evaluation Details
E.1 Bias Score Metrics in BBQ
@@ E.2 Prompt for Scoring on CEB
Appendix F Full Results for Llama2-13B on BBQ
@@
Appendix G Additional Case Studies